OpenStudy (anonymous):

Solve, simplify answer. 1/2log(base4)x=2log(base4)2

OpenStudy (saifoo.khan):

\[\log_4 x^{0.5} = \log_4 2^2\]

OpenStudy (saifoo.khan):

\[x^{0.5} = 2^2\] \[\sqrt x = 4\]take square on both sides, \[(\sqrt x)^2 = (4)^2\] \[x = 16\]

OpenStudy (anonymous):

oh ok...thank you

OpenStudy (saifoo.khan):

no problem.

OpenStudy (anonymous):

gt